### v2

# Circle bundles with PSC over some four manifolds

Abstract.
We construct infinitely many examples of four manifolds with macroscopic dimension $4$ equipped with circle bundles whose total spaces admit metrics of positive scalar curvature. Further, we verify that these bundles have macroscopic dimension at most $3$. In particular, we answer a question of Gromov on the existence of circle bundles over enlargeable manifolds whose total spaces admit metrics of positive scalar curvature. Our constructions are based on techniques from symplectic geometry.
The second author is supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500.
1. Introduction
1.1. Background and motivation
Definition 1 (Urysohn width).
Let $X$ be a metric space. The $k$-Urysohn width of $X$, denoted by $\mathrm{UW}_{k}(X)$, is the infimum over $t>0$ such that there is a $k$-dimensional simplicial complex $Y$ and a continuous map $f:X\to Y^{k}$ satisfying the property that the diameter of each fiber of $f$ is at most $t$, i.e. for all $y\in Y$, we have $\mathrm{diam}(f^{-1}\{y\})\leq t$.
If $k<n$, then a small $k$-Urysohn width intuitively means that the manifold is close to a $k$-dimensional space. This notion of closeness is captured in the following definition.
Definition 2 (Macroscopic dimension).
Let $X$ be a metric space. The macroscopic dimension, denoted by $\dim_{mc}X$, is the minimal $k$ such that $\mathrm{UW}_{k}(X)<\infty$. In other words, $\dim_{mc}X\leq k$ if there is a continuous map $f:X\to Y^{k}$ to a $k$-dimensional simplicial complex $Y$, such that for all $y\in Y$, $\mathrm{diam}f^{-1}(y)<C$ for some uniform constant $C$.
Remark.
Note that for a compact metric space $X$, $\dim_{mc}(X)=0$. Therefore, when one refers to the macroscopic dimension of a closed manifold $M$, one typically means the macroscopic dimension of the universal cover $\widetilde{M}$.
The notions of Urysohn width and macroscopic dimension appear in several works of Gromov, with one of the aims being to capture the appropriate notion of smallness that must be exhibited by a manifold admitting a metric of positive scalar curvature (PSC). These also appear naturally in the study of non-collapsing and thick-thin decomposition. See also some of these recent works on the topic [ABG24, Gut11, Nab22, Sab22]. The primary motivation behind this article is the following conjecture formulated by Gromov for manifolds with positive scalar curvature [Gro96, Section $2\text{\textonehalf}$].
Conjecture 1 (Gromov).
Let $M$ be a closed connected $n$-dimensional manifold, admitting a PSC metric. Then the macroscopic dimension of the universal cover of $M$ is at most $n-2$, that is,
$\dim_{mc}\widetilde{M}\leq n-2.$
Intuitively, the conjecture says that a manifold admitting a PSC metric, on large scales, looks like a manifold of at least two dimensions less. The primary evidence in this direction is the following simple observation: Given any closed manifold $M$, consider the manifold $M\times S^{2}_{r}$ with the product metric. Here, $S^{2}_{r}$ denotes a sphere of radius $r$. Then for $r$ small enough, the product metric on $M\times S^{2}_{r}$ has positive scalar curvature. In light of the previous remark, we pass to the universal cover of $M\times S^{2}_{r}$, which is $\widetilde{M}\times S^{2}_{r}$. Since the projection map to $\widetilde{M}$ is continuous and each fiber is $S^{2}_{r}$, we have
$\mathrm{UW}_{n-2}(\widetilde{M\times S^{2}_{r}})\leq 2r,$
and therefore the macroscopic dimension $\dim_{mc}(\widetilde{M\times S^{2}_{r}})\leq n-2$. Conjecture 1 asserts that all PSC manifolds must exhibit this behavior. Gromov’s conjecture has been verified in dimension $3$ [GL83, LM23]. In higher dimensions there are several partial results under various assumptions (cf. [BD16, DD22]).
While Conjecture 1 arises from the simple observation that existence of a PSC metric is guaranteed given the presence of an $S^{2}$ factor in the manifold, it is also natural to ask about the possible interaction between existence of PSC metrics and having an $S^{1}$ factor. This has been widely investigated. For instance, PSC manifolds with a trivial $S^{1}$ factor (i.e., $M\times S^{1}$) are amenable to the Schoen-Yau conformal descent argument [SY79], as they have a non-trivial codimension-$1$ homology class given by the Poincaré dual of the $S^{1}$ factor. Such $S^{1}$ factors are also useful in quantitative geometric problems related to scalar curvature. For example, building on an observation of Fisher-Colbrie and Schoen [FCS80], Gromov and Lawson introduced the toric symmetrization procedure [GL83] which was used to estimate the Lipschitz constant of a non-zero degree map from a PSC manifold. More recently, Gromov [Gro18] observed that toric symmetrization can also be used to prove width estimates for Riemannian bands. On the same theme, Rosenberg [Ros07, Conjecture 1.24] formulated the following conjecture regarding the relationship between PSC metrics and trivial $S^{1}$ factors.
Conjecture 2 (Rosenberg).
Let $M$ be a closed connected $n$-dimensional manifold. Then $M\times S^{1}$ admits a PSC metric iff $M$ admits a PSC metric.
However, the situation is less clear when the $S^{1}$ factor arises as a fiber of a non-trivial circle bundle. The key difficulty in studying existence of PSC metrics on non-trivial circle bundles is that, in this case, the Schoen-Yau conformal descent argument does not apply in a straightforward way. [^1] Indeed, the integral homology class of the $S^{1}$ fiber in nontrivial circle bundles is always torsion, with order given by the divisibility of the Euler class of the bundle. Therefore, its Poincaré dual is zero. In this direction, Gromov posed the following question [Gro18, pg. 661]:
Question 1.
[Gro18, pg. 661] “What happens to nontrivial bundles $\underline{X}\to Y$?
[…] In general, the examples in [BH10] indicate a possibility of non-enlargeable circle bundles over enlargeable $Y$; yet, it seems hard(er) to find such examples, where the corresponding $\underline{X}$ would admit complete metrics with $Sc\geq\sigma>0$.”
In the negative direction we point out the recent comprehensive work [He23] on PSC metrics on non-trivial circle bundles due to He. In particular, he showed that circle bundles with homologically non-trivial fibers do not admit PSC metrics if the base space is a $1$-enlargeable2221111-enlargeable means that the maps to $S^{n}$ in the definition of enlargeability are of degree $1$. manifold of dimension $n\leq 6$ [He23, Proposition 4.12].
The main goal of this article is to study and exhibit examples of interactions between the existence of a positive scalar curvature metric and macroscopic dimension on non-trivial circle bundles over some $4$-manifolds. In particular, we give a positive answer to Question 1. We construct these examples using techniques from symplectic geometry. The inspiration for our construction came from the following remark by Gromov [Gro23, pg. 8]:
The emergent picture of spaces with $Sc.curv\geq 0$, where topology and geometry are intimately intertwined, is reminiscent of the symplectic geometry, but the former has not reached yet the maturity of the latter.
Granting this dictum of similarity between scalar curvature geometry and symplectic geometry, one is led to believe that the analogue of the Schoen-Yau descent for positive scalar curvature manifolds, in the symplectic case must be Donaldson’s symplectic divisor theorem [Don96], as it can be viewed as a codimension-$2$ descent-type result for symplectic manifolds. This analogy is strengthened when seen in conjunction with a recent result of Cecchini, Räde and Zeidler [CRZ23] which establishes a certain codimension-$2$ descent for positive scalar curvature manifolds of dimension $7$. In fact, we show that circle bundles over Donaldson divisors in symplecic $6$-manifolds fit neatly into the framework of applicability of [CRZ23].
1.2. Main results
We discuss the necessary symplectic preliminaries in Section 2.
In this article, we construct infinitely many examples of $4$-manifolds $Z$ with arbitrarily large Euler number $|\chi(Z)|$, arising as Donaldson divisors (see Theorem 1) in symplectic $6$-manifolds, such that $Z$ does not admit a PSC metric and macroscopic dimension of $\widetilde{Z}$ is exactly $4$. Nevertheless, we construct circle bundles over these $Z$,
[…]
Note that this phenomenon of the macroscopic dimension decreasing upon taking a (twisted) product with $S^{1}$ is not possible in the case of a trivial product. Indeed,
$\dim_{mc}(\widetilde{Z\times S^{1}})=\dim_{mc}(\widetilde{Z}\times\mathbf{R})=\dim_{mc}(\widetilde{Z})+1.$
To the best of our knowledge examples exhibiting a drop in macroscopic dimension on twisting with $S^{1}$ have not appeared before in literature. On a similar theme, we were informed by Guth [Gut24] that a drop in the Urysohn width upon taking a double cover, i.e. an $S^{0}$ bundle, was observed by Balitskiy [Bal21, Theorem 2.4.4, pg. 36].
We now state the main result of this article:
Theorem A.
[…]
  1. (1)
    The total space $E_{Z}$ admits a metric of positive scalar curvature.
  2. (2)
    The macroscopic dimension of the total space $E_{Z}$ is at most 3.
  3. (3)
    If $b_{2}^{+}(M)>1$, then $Z$ does not admit a positive scalar curvature metric.
  4. (4)
    If $M$ admits a positive degree map to an aspherical manifold, then $Z$ does not admit a positive scalar curvature metric.
  5. (5)
    If $M$ is a symplectically aspherical [^4] See Definition 3. In particular, a symplectic manifold with $\pi_{2}=0$ (e.g. $T^{4}$) is symplectically aspherical. manifold and $\pi_{1}(M)$ is amenable, then the macroscopic dimension of $Z$ is exactly $4$.
We emphasize that Conclusions $(1)$ and $(2)$ make no assumptions either on the fundamental groups of $M$ and $E_{Z}$, or on the (non-)existence of spin or almost spin structures on the universal cover of $E_{Z}$. In particular, Conclusion $(2)$ does not follow from any existing results on Gromov’s conjecture (Conjecture 1). On the other hand, Conclusions $(3)$, $(4)$ and $(5)$ rest on Lemma 5, where we will show that there is a positive degree map from $Z$ to $M$. Then Conclusion $(3)$ follows from a theorem of Taubes [Tau94], Conclusion $(4)$ follows from the work of Chodosh, Li and Liokumovich [CLL23] and Conclusion $(5)$ follows from combining Proposition 6 with a result of Dranishnikov [Dra11].
By combining Theorem A with a Euler characteristic computation for Donaldson divisors carried out in Proposition 4, we obtain the following uniqueness corollary:
Corollary B.
There exists infinitely many $4$-manifolds $Z_{k}$ and circle bundles $E_{k}$ over $Z_{k}$ such that $|\chi(Z_{k})|\sim O(k^{3})$, and
  1. (1)
    $\dim_{mc}Z_{k}=4$ but $\dim_{mc}E_{k}=3$,
  2. (2)
    $Z_{k}$ is not PSC and $E_{k}$ is PSC.
As a further corollary of Theorem A, we will construct examples sought after in Gromov’s Question 1.
Corollary C.
There exists enlargeable closed $4$-manifolds $Z$ and nontrivial circle bundles $E\to Z$ such that $E$ admits a positive scalar curvature metric.
1.3. Organisation
In Section 2, we define the necessary symplectic preliminaries and make some observations regarding Donaldson divisors. In Section 3, we assemble other lemmas and propositions that we need to prove our main results. In Section 4, we prove Theorem A and Corollary C as well as provide explicit examples that are used towards proofs of Corollary B. The reader may first read Section 4, and refer to Sections 2 and 3, for the proofs of the results used in Section 4 as they arise.
Acknowledgements
The authors would like to thank Shihang He for several helpful comments on the first version of this paper. They would also like to thank Alexey Balitskiy, Alexander Dranishnikov, and Larry Guth for answering their questions and providing helpful remarks. The first author wishes to thank his advisor Yannick Sire for support and encouragement. The second author wishes to thank his advisor Mahan Mj, as well as his friends and colleagues Ritwik Chakraborty, and Sekh Kiran Ajij for several helpful conversations.
2. Donaldson divisors: definition and properties
Let $(X,\omega)$ be a symplectic manifold. The symplectic form $\omega$ is said to be integral if the cohomology class $[\omega]\in H^{2}(X;\mathbf{R})$ lies in the image of the change of coefficients homomorphism $H^{2}(X;\mathbf{Z})\to H^{2}(X;\mathbf{R})$.
Remark.
Any symplectic manifold $(X,\eta)$ admits a symplectic form $\omega$ which is integral. To see this, choose a collection of closed $2$-forms on $X$ representing a $\mathbf{Q}$-basis of the vector space $H^{2}(X;\mathbf{Q})\subset H^{2}(X;\mathbf{R})$, and expresses the class $[\eta]\in H^{2}(X;\mathbf{R})$ as a $\mathbf{R}$-linear combination of the chosen basis elements. Approximate the real coefficients by rational numbers and then multiply by an integer to clear out the denominators. Let $\omega$ be the resulting closed $2$-form. Then $\omega$ is an integral symplectic form on $X$. Indeed, one only needs to verify that $\omega$ is non-degenerate. But it is so, since non-degeneracy is both an open condition and is scale-invariant.
The following foundational result in symplectic geometry was established by Donaldson [Don96].
Theorem 1.
[Don96, Theorem 1] Suppose $(X^{2n},\omega)$ is a symplectic manifold where $\omega$ is integral. Then, for sufficiently large $k\gg 1$, there exists a symplectic submanifold $Z\subset X$ of codimension $2$, such that $Z$ represents the class Poincaré dual to $k[\omega]$.
The symplectic submanifolds $Z\subset(X,\omega)$ appearing in the statement of the theorem are referred to as Donaldson divisors of $(X,\omega)$. We briefly outline the construction of $Z$. Since $\omega$ is an integral symplectic form, we can consider the pre-quantum line bundle
$\pi:L\to X$
for $(X,\omega)$. This is a complex line bundle with a hermitian connection $\nabla$ such that the curvature $2$-form is given by $F_{\nabla}=-2\pi i\omega$. Thus, note that the first Chern class of the line bundle $L$ is $c_{1}(L)=[\omega]$. The symplectic submanifold $Z^{2n-2}\subset(X^{2n},\omega)$ is produced as the zero set of an asymptotically holomorphic section of $L^{\otimes k}$. For details, see [Don96, Theorem 5]. Note $c_{1}(L^{\otimes k})=k[\omega]$. Since $Z$ is the zero set of a section of $L^{\otimes k}$, we must have that $Z$ represents the class Poincaré dual to $k[\omega]$.
2.1. Complement of Donaldson divisors
We note the following two simple but useful observations.
Lemma 2.1.
Let $L^{\otimes k}|_{Z}$ denote the restriction of $L^{\otimes k}$ to $Z$. Denote by $\mathbb{S}(L^{\otimes k}|_{Z})$ the unit circle bundle of $L^{\otimes k}|_{Z}$. Then the following statements are true.
  1. (1)
    The tubular neighborhood $\nu(Z)$ of $Z\subset X$ is diffeomorphic to $L^{\otimes k}|_{Z}$
  2. (2)
    The boundary $\partial\nu(Z)$ is diffeomorphic to $\mathbb{S}(L^{\otimes k}|_{Z})$.
Proof.
By the tubular neighborhood theorem, $\nu(Z)$ is diffeomorphic to the normal bundle of $Z\subset X$. By the construction of $Z$ outlined above, $Z=s^{-1}(\mathbf{0}),$ where $s:X\to L^{\otimes k}$ is a section transverse to the zero section $\mathbf{0}\subset L^{\otimes k}$. Since the normal bundle to $\mathbf{0}\subset L^{\otimes k}$ is $L^{\otimes k}$ itself, the normal bundle to $Z$ in $X$ is $L^{\otimes k}|_{Z}$ by transversality. Since $\nu(Z)\cong L^{\otimes k}_{Z}$, the total spaces of the unit circle bundles are also diffeomorphic, i.e. $\partial\nu(Z)\cong\mathbb{S}(L^{\otimes k}|_{Z})$. ∎
The following is an analogue of the Lefschetz hyperplane theorem in the symplectic setting, which we shall use in the next section:
Proposition 2.
[Don96, Proposition 39] The inclusion $j:Z^{2n-2}\hookrightarrow X^{2n}$ induces an isomorphism $j_{*}:\pi_{k}(Z)\to\pi_{k}(X),$ for all $k\leq n-2$, and a surjection for $k=n-1$.
In fact, the following lemma can be deduced from the proof of Proposition 2 in [Don96, pg. 700-701]:
Lemma 2.2.
The complement $X^{2n}\setminus Z$ deformation retracts to a simplicial complex of dimension $n$.
Proof.
By the construction of $Z\subset X$ summarized above, there is an asymptotically holomorphic section $s:X\to L^{\otimes k}$ such that $Z=s^{-1}(\mathbf{0})$. Let $\psi:X\setminus Z\to\mathbf{R}$ be defined by $\psi:=-\log|s|^{2}$. During the course of the proof of [Don96, Proposition 39], the author shows that $\psi$ is a Morse function on $X\setminus Z$ with critical points having index at most $n$. Therefore, by Morse theory, $X\setminus Z$ deformation retracts to an $n$-dimensional simplicial complex. ∎
We deduce two corollaries from Lemma 2.2. These are central to the proof of Part (2) of Theorem A.
Corollary 3.
Let $\nu(Z)$ be a tubular neighborhood of $Z$ in $X^{2n}$. Let $W=X\setminus\nu(Z)$. Then the following statements are true,
  1. (1)
    $W$ deformation retracts to a simplicial complex of dimension $n$.
  2. (2)
    If $n\geq 3$, the inclusion $\partial W\hookrightarrow W$ induces an isomorphism in $\pi_{1}$.
Proof.
For Part (1), observe $\partial W=\partial\nu(Z)$. So, by Lemma 2.1, we have $\partial W\cong\mathbb{S}(L^{\otimes k}|_{Z})$. Therefore, $X\setminus Z$ is simply $W$ with a collar $\partial W\times[0,\infty)$ attached. Consequently, $X\setminus Z$ deformation retracts to $W$. Since by Lemma 2.2, $X\setminus Z$ deformation retracts to a $n$-dimensional simplicial complex, so does $W$.
For Part $(2)$, observe that the Morse function $\psi$ constructed in Lemma 2.2 gives a handlebody decomposition of $W$ by handles of index at most $n$. By reversing the handlebody decomposition, we get that $W$ can be constructed from $\partial W$ by attaching handles of index at least $n$. If $n\geq 3$, then we conclude that all these handles are of index at least $3$. Since $\pi_{1}$ is unchanged under attachements of handles of index bigger than $2$, we conclude $\partial W\hookrightarrow W$ induces an isomorphism in $\pi_{1}$. ∎
2.2. Topology of Donaldson divisors in dimension $6$
In the next proposition, we show that Donaldson divisors of a $6$-dimensional symplectic manifold $(X^{6},\omega)$ with integral symplectic form $\omega$, realizing the homology classes Poincaré dual to $k[\omega]$, are topologically distinct for distinct values of $k\gg 1$. We do this by calculating their Euler characteristic.
Proposition 4.
Let $Z_{k}\subset(X^{6},\omega)$ denote a Donaldson divisor with $[Z_{k}]=\mathrm{PD}(k[\omega])$. Then, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|\sim O(k^{3})$ is monotonically increasing in $k$. In particular, for $k,l\gg 1$ sufficiently large, $Z_{k}\cong Z_{l}$ if and only if $k=l$.
Proof.
From Lemma 2.1, we know that the normal bundle of $Z_{k}$ in $X$ is isomorphic to $L^{\otimes k}|_{Z_{k}}$, where $L$ is a complex line bundle over $X$ with $c_{1}(L)=[\omega]$. Therefore,
$TZ_{k}\oplus L^{\otimes k}|_{Z_{k}}\cong TX|_{Z_{k}}$
Let $j:Z_{k}\hookrightarrow X$ denote the inclusion map. We compute the Chern classes:
$c_{1}(TZ_{k})+c_{1}(L^{\otimes k}|_{Z_{k}})=j^{*}c_{1}(TX)$ (2.1)
$c_{2}(TZ_{k})+c_{1}(TZ_{k})\smile c_{1}(L^{\otimes k}|_{Z_{k}})=j^{*}c_{2}(TX)$ (2.2)
Note that $c_{1}(L^{\otimes k}|_{Z_{k}})=j^{*}c_{1}(L^{\otimes k})=k\cdot j^{*}[\omega]$. Substituting this in (2.1), we obtain:
$c_{1}(TZ_{k})=j^{*}c_{1}(TX)-k\cdot j^{*}[\omega]$ (2.3)
Substituting (2.3) in (2.2), we obtain
$c_{2}(TZ_{k})$ $=j^{*}c_{2}(TX)-j^{*}c_{1}(TX)\smile c_{1}(L^{\otimes k}|_{Z_{k}})+k\cdot j^{*}[\omega]\smile c_{1}(L^{\otimes k}|_{Z_{k}})$
$=j^{*}c_{2}(TX)-k\cdot j^{*}(c_{1}(TX)\smile[\omega])+k^{2}\cdot j^{*}[\omega]^{\smile 2}$ (2.4)
Since the top Chern class is the Euler class, evaluating the cohomology classes appearing in either sides of (2.4) on $[Z_{k}]=k\cdot\mathrm{PD}[\omega]$ we get:
$\chi(Z_{k})$ $=\int_{Z_{k}}c_{2}(TZ_{k})$
$=k^{3}\int_{M}[\omega]^{\smile 3}-k^{2}\int_{M}c_{1}(TX)\smile[\omega]^{\smile 2}+k\int_{M}c_{2}(TX)\smile[\omega]$
Note that $c_{1}(TX),c_{2}(TX),[\omega]$ are all independent of $k$. Hence, $\chi(Z_{k})$ is a cubic polynomial in $k$ with coefficients independent of $k$, and a nonzero leading coefficient. Therefore, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|$ is monotonically increasing in $k$, as required. ∎
Remark.
The fact that $\chi(Z_{k})\sim O(k^{3})$ in $X^{6}$ should be put in context of the following standard observation: Let $X=\mathbf{CP}^{3}$ and $V_{k}\subset\mathbf{CP}^{3}$ be a degree $k$ complex hypersurface, then $\chi(V_{k})=k^{3}-4k^{2}+6k$.
2.3. Donaldson divisors of $M\times S^{2}$
Let $(M^{4},\eta_{0})$ be an integral symplectic $4$-manifold. We equip the manifold $M\times S^{2}$ with the integral symplectic form $\omega=\eta_{0}\oplus\omega_{0}$, where $\omega_{0}$ denotes the standard area form on $S^{2}$. We now make the following observation, which will be crucial to the proof of Parts (4) and (5) of Theorem A.
Proposition 5.
The map $f:Z\to M$, obtained by the composition
[…]
has positive degree.
Proof.
Take any regular value $p\in M$. Then the oriented count #$\#f^{-1}(p)$ is equal to the oriented intersection number #$\#([Z]\cap[\{p\}\times S^{2}])$. Therefore, we have
$\deg(f)$ =#$=\#([Z]\cap[\{p\}\times S^{2}])$
[…]
$>0$
The last inequality holds because $\{p\}\times S^{2}$ is a symplectic submanifold of $(M\times S^{2},\omega)$. ∎
2.4. Symplectically aspherical manifolds
Definition 3.
A symplectic manifold $(X,\omega)$ is said to be symplectically aspherical if for every smooth map $\phi:S^{2}\to X$,
$\int_{S^{2}}\phi^{*}\omega=0$
Remark.
Note that it is immediate from Definition 3 that symplectic submanifolds of symplectically aspherical manifolds are also symplectically aspherical. Since symplectic manifolds $(X,\omega)$ with $\pi_{2}(X)=0$ are automatically symplectically aspherical, together with the previous observation as well as Donaldson’s theorem (Theorem 1), this provides a large class of symplectically aspherical manifolds.
This class of manifolds was originally introduced by Floer in the context of the Arnol’d conjecture in symplectic geometry. We refer the reader to the survery article [KRT08] for detailed information. The key property of symplectically aspherical manifolds that we shall be using is the following:
Proposition 6.
Let $(X^{2n},\omega)$ be an integral symplectic manifold which is symplectically aspherical. Then $X$ is a rationally essential manifold. [^5] A $d$-dimensional manifold $M^{d}$ is rationally essential if for the map $\phi:M\to K(\pi_{1}(M),1)$ classifying the universal cover of $M$, $\phi_{*}[M]\neq 0\in H_{d}(K(\pi_{1}(M),1);\mathbf{Q})$
Proof.
By attaching cells of dimension $\geq 3$ to $X$, we kill the higher homotopy groups $\pi_{i}(X)$, $i\geq 2$ to obtain a CW-complex which is a model for a $K(\pi_{1}(X),1)$. By construction, there is a natural inclusion as a subcomplex
$i:X\hookrightarrow K(\pi_{1}(X),1)$
Let $L$ be the pre-quantum line bundle over $X$, with $c_{1}(L)=[\omega]$. Since $(X,\omega)$ is symplectically aspherical, $L$ must extend to the $3$-skeleton of $K(\pi_{1}(X),1)$. Indeed, for any $3$-cell of $K(\pi_{1}(X),1)$, let $\phi:S^{2}\to X$ denote the attaching map. Then, we have:
$\langle c_{1}(\phi^{*}L),[S^{2}]\rangle=\int_{S^{2}}\phi^{*}\omega=0$
Therefore, $c_{1}(\phi^{*}L)=0$. Consequently, $\phi^{*}L$ must be the trivial bundle over the $2$-sphere. Thus, we may extend $L$ as a trivial bundle over the $3$-cell. Since complex line bundles over higher dimensional spheres are always trivial, we can extend $L$ over all the other higher dimension cells as well.
In this way, we obtain a line bundle $L^{\prime}$ over $K(\pi_{1}(X),1)$ such that $i^{*}L^{\prime}=L$. Let us denote $\alpha:=c_{1}(L^{\prime})\in H^{2}(K(\pi_{1},1);\mathbf{Z})$. Then, $i^{*}\alpha=[\omega]$. Consequently, $i^{*}\alpha^{\wedge n}=[\omega^{\wedge n}]\neq 0$. Thus, $i^{*}:H^{2n}(K(\pi_{1}(X),1);\mathbf{Z})\to H^{2n}(X;\mathbf{Z})$ is nonzero. Since $[\omega^{\wedge n}]$ is not torsion, $\alpha^{\wedge n}$ must not be torsion either. Therefore, $i_{*}:H^{2n}(K(\pi_{1},1);\mathbf{Q})\to H^{2n}(X;\mathbf{Q})$ is nonzero. Therefore, by the universal coefficients theorem, $i_{*}:H_{2n}(X;\mathbf{Q})\to H_{2n}(K(\pi_{1}(X),1);\mathbf{Q})$ must be nonzero as well. ∎
3. Preliminary results
In this section we assemble all other results that we will require to prove Theorem A.
3.1. Nullcobordism and macroscopic dimension
We first prove a general lemma relating macroscopic dimension of a manifold with the homotopy type of a nullcobordism of the manifold. This will be used to prove of Part (2) of Theorem A.
Lemma 3.1.
Let $W^{n}$ be a compact $n$-dimensional manifold with boundary $E=\partial W$. Suppose
  1. (1)
    $W$ deformation retracts to a $k$-dimensional simplicial complex $K\subset W$,
  2. (2)
    The inclusion $i:E=\partial W\hookrightarrow W$ induces an isomorphism in $\pi_{1}$.
Then, $\dim_{mc}\widetilde{E}\leq k$.
Proof.
Let $r:W\to K$ denote the deformation retract. Let $f=r\circ i:E\to K$ denote the composition with the inclusion $i:E\hookrightarrow W$. Since $i$ and $r$ both induce an isomorphism in $\pi_{1}$, so does $f$. Therefore, $f$ lifts to a map between the respective universal covers:
$\widetilde{f}:\widetilde{E}\to\widetilde{K}$
Choose an arbitrary Riemannian metric $g$ on $W$, and let $\widetilde{g}$ be its lift to the universal cover $\widetilde{W}$. We will show that the fibers of $\widetilde{f}$ are uniformly bounded with respect to the metric $\widetilde{g}$. This will give us the desired conclusion, i.e. $\dim_{mc}\widetilde{E}\leq k$.
To this end, let us assume without loss of generality that $f$ is a simplicial map. This can be arranged by the simplicial approximation theorem. Let $\widetilde{p}\in\widetilde{K}$ be a point lying over $p\in K$. Then, we claim $\widetilde{f}^{-1}(\widetilde{p})$ is isometric to $f^{-1}(p)$. Indeed, note that $f^{-1}(p)\hookrightarrow M$ induces the zero map on $\pi_{1}$, since otherwise there would be a loop $\gamma\subset f^{-1}(p)$ which is not nullhomotopic in $M$ such that $f_{*}[\gamma]=0\in\pi_{1}(K)$. This would be absurd, as $f$ induces an isomorphism on $\pi_{1}$. Thus, $f^{-1}(p)$ lifts homeomorphically in the cover to a unique subset in $\widetilde{Y}$ containing $\widetilde{p}$. By uniqueness, this lift is exactly $\widetilde{f}^{-1}(\widetilde{p})$. Consequently, the covering projection must be a homeomorphism:
$\widetilde{f}^{-1}(\widetilde{p})\to f^{-1}(p)$
Since the metric on the cover is $\widetilde{g}$, obtained by lifting the metric $g$ on $W$, the aforementioned homeomorphism is the desired isometry. Thus, the fibers of $\widetilde{f}$ are isometric to the fibers of $f$, which are uniformly bounded as $E$ is compact. Therefore, $\dim_{mc}\widetilde{E}\leq k$. ∎
3.2. Positive scalar curvature and Donaldson divisors
We begin with the following result due to Cecchini, Räde and Zeidler in [CRZ23], which establishes a codimension-$2$ obstruction to existence of a PSC metric on a $7$-manifold.
Theorem 2.
[CRZ23, Theorem 1.11] Let $Y^{7}$ be a closed connected manifold, and $X^{5}\subset Y^{7}$ be a submanifold such that
  1. (1)
    $i:X\hookrightarrow Y$ has trivial normal bundle,
  2. (2)
    $i_{*}:\pi_{1}(X)\to\pi_{1}(Y)$ is injective,
  3. (3)
    $i_{*}:\pi_{2}(X)\to\pi_{2}(Y)$ is surjective
Then, if $X$ does not admit a PSC metric, $Y$ does not admit a PSC metric either.
Let $(M^{6},\omega)$ be an integral symplectic $6$-manifold. Let $Z\subset(M^{6},\omega)$ be a Donaldson divisor. Recall from Lemma 2.1 that the normal bundle of $Z$ in $M$ is isomorphic to $L^{\otimes k}$, which is a nontrivial complex line bundle over $Z$. However, we will now consider the unit circle bundle of $L^{\otimes k}$, denoted by $\mathbb{S}(L^{\otimes k})$. In conjunction with it, we will consider the restriction of the unit circle bundle of $L^{\otimes k}$ to $Z$, denoted by $\mathbb{S}(L^{\otimes k}|_{Z})$. Note that $\mathbb{S}(L^{\otimes k})$ is a $7$-manifold and $\mathbb{S}(L^{\otimes k}|_{Z})$ is a $5$-manifold.
In the following proposition, we shall verify that the pair $(\mathbb{S}(L^{\otimes k}),\mathbb{S}(L^{\otimes k}|_{Z}))$ satisfies the hypotheses of Theorem 2.
Proposition 7.
Let us denote $Y:=\mathbb{S}(L^{\otimes k})$ and $X:=\mathbb{S}(L^{\otimes k}|_{Z})$. Then the pair $(Y^{7},X^{5})$ satisfies the hypotheses of Theorem 2. Therefore, if $\mathbb{S}(L^{\otimes k})$ is PSC, then so is $\mathbb{S}(L^{\otimes k}|_{Z})$.
Before the proof of Proposition 7, we state and prove an elementary topology lemma.
Lemma 3.2.
Let $\pi:E\to B$ be a complex line bundle, and $\mathbb{S}(E)\subset E$ be the unit circle bundle with respect to some choice of a fiberwise Riemannian metric on $E$. Let $\pi_{s}:\mathbb{S}(E)\to B$ denote the bundle projection. Then, $\pi_{s}^{*}E$ is trivial over $\mathbb{S}(E)$.
Proof.
By definition of the pullback construction,
$\pi_{s}^{*}E=\mathbb{S}(E)\times_{B}E$
The diagonal section $\sigma:\mathbb{S}(E)\to\mathbb{S}(E)\times_{B}E$ defined by $\sigma(e):=(e,e)$ is a nowhere zero section of the circle bundle $\pi_{s}^{*}E\to\mathbb{S}(E)$. Therefore, $\pi_{s}^{*}E$ is trivial over $\mathbb{S}(E)$. ∎
Proof of Proposition 7.
First, we observe that the normal bundle of $X\subset Y$ is trivial. Indeed, let $E:=L^{\otimes k}|_{Z}$ and $\pi_{s}:\mathbb{S}(E)=X\to Z$ denote the bundle projection. Then,
$N_{Y}X=\pi^{*}N_{M}Z\cong\pi_{s}^{*}(E)$
By Lemma 3.2, we conclude $N_{Y}X$ is trivial over $\mathbb{S}(E)=X$, as desired.
Next, we show $i$ induces an isomorphism on $\pi_{1}$ and a surjection on $\pi_{2}$. To this end, consider the following commutative diagram:
[…]
The rows are fiber bundles, therefore by applying the long exact sequence in homotopy and using naturality, we have the following diagram:
00{0}${{\pi_{2}(Y)}}$${{\pi_{2}(M)}}$${{\pi_{1}(S^{1})\cong\mathbf{Z}}}$${{\pi_{1}(Y)}}$${{\pi_{1}(M)}}$00{0}00{0}${{\pi_{2}(X)}}$${{\pi_{2}(Z)}}$${{\pi_{1}(S^{1})\cong\mathbf{Z}}}$${{\pi_{1}(X)}}$${{\pi_{1}(Z)}}$00{0}$\scriptstyle{\phi_{M}}$$\scriptstyle{i_{*}}$$\scriptstyle{j_{*}}$$\scriptstyle{\phi_{Z}}$$\scriptstyle{\mathrm{id}}$$\scriptstyle{i_{*}}$$\scriptstyle{j_{*}}$
Note, as indicated in the diagram above, $j_{*}:\pi_{1}(Z)\to\pi_{1}(M)$ is an isomorphism and $j_{*}:\pi_{2}(Z)\to\pi_{2}(M)$ is a surjection due to Proposition 2 specialized to the case of $n=3$, as $M$ is a symplectic manifold of dimension $62n=$ 6.
  1. To show $i_{*}:\pi_{2}(X)\to\pi_{2}(Y)$ is surjective:
    Let $\alpha\in\pi_{2}(Y)$. Let $\alpha^{\prime}\in\pi_{2}(M)$ be the image of $\alpha$. Then $\phi_{M}(\alpha^{\prime})=0$. Since $j_{*}:\pi_{2}(Z)\to\pi_{2}(M)$ is surjective, there exists $\beta^{\prime}\in\pi_{2}(M)$ such that $j_{*}(\beta^{\prime})=\alpha^{\prime}$. By commutativity,
    $\phi_{Z}(\beta^{\prime})=\phi_{M}(j_{*}(\beta^{\prime}))=\phi_{M}(\alpha^{\prime})=0.$
    Therefore, $\beta^{\prime}$ lies in the kernel of $\phi_{Z}$, which is equal to the image of $\pi_{2}(X)$ in $\pi_{2}(Z)$. Thus, there exists $\beta\in\pi_{2}(X)$ such that image of $\beta$ is exactly $\beta^{\prime}$. Now, the image of $i_{*}(\beta)-\alpha$ in $\pi_{2}(M)$ is $j_{*}(\beta^{\prime})-\alpha^{\prime}=0$. Since $\pi_{2}(Y)$ injects into $\pi_{2}(M)$, we conclude $i_{*}(\beta)=\alpha$. This proves the claim.
  2. To show $i_{*}:\pi_{1}(X)\to\pi_{1}(Y)$ is injective:
    Note that $\phi_{M}$ and $\phi_{Z}$ have the same image and therefore the same cokernel. By the short five-lemma, $i_{*}:\pi_{1}(X)\to\pi_{1}(Y)$ is an isomorphism. In particular, it is injective. ∎
4. Proofs of the main results
4.1. Proof of Theorem A
In this section, we prove our main result. Let $(M,\eta)$ be an integral symplectic $4$-manifold. Let $\omega_{0}$ be the standard area form on $S^{2}$. We equip $M\times S^{2}$ with the integral symplectic form $\omega=\eta\oplus\omega_{0}$. Let $L$ be the pre-quantum line bundle on the symplectic manifold $(M\times S^{2},\omega)$. Let $Z\subset(M\times S^{2},\omega)$ be a Donaldson divisor, with $[Z]=\mathrm{PD}(k\cdot[\omega])$, for $k\gg 1$ sufficiently large. We set $E_{Z}:=\mathbb{S}(L^{\otimes k}|_{Z})$ as the unit circle bundle of $L^{\otimes k}$ restricted to $Z$ (as introduced in Lemma 2.1 and Proposition 7).
  1. Proof of (1).
    As $M\times S^{2}$ admits a PSC metric, and $\mathbb{S}(L^{\otimes k})$ is a circle bundle on $M\times S^{2}$, $\mathbb{S}(L^{\otimes k})$ also admits a PSC metric [BB83, Theorem C]. By Proposition 7, this implies that $E_{Z}=\mathbb{S}(L^{\otimes k}|_{Z})$ also admits a PSC metric.
  2. Proof of (2).
    Let $\nu(Z)$ denote a tubular neighbourhood of $Z\subset M\times S^{2}$. Consider the manifold with boundary $W:=(M\times S^{2})\setminus\nu(Z)$. Note that $\partial W=\partial\nu(Z)$. By Lemma 2.1, $\partial\nu(Z)\cong\mathbb{S}(L^{\otimes k}|_{Z})=E_{Z}$. Therefore $\partial W\cong E_{Z}$. In Corollary 3, we deduced that $(W^{6},E_{Z})$ satisfies the hypothesis of Lemma 3.1 with $n=6$ and $k=3$. Therefore, $\dim_{mc}\widetilde{E_{Z}}\leq 3$.
  3. Proof of (3).
    By Proposition 5, the map $f:Z\to M$ is of positive degree. Moreover, by hypothesis $b_{2}^{+}(M)>1$. Since $f$ is of positive degree, the induced map
    $f^{*}:H^{2}(M;\mathbf{Q})\to H^{2}(Z;\mathbf{Q})$
    is an injective map of quadratic spaces, where both domain and range are equipped with the quadratic form given by the cup product $\smile$. Therefore, $b_{2}^{+}(Z)>1$. In [Tau94], Taubes proved that symplectic $4$-manifolds with $b_{2}^{+}>1$ have non-zero Seiberg-Witten invariant. Since having non-zero Seiberg-Witten invariant obstructs the existence of PSC metrics, this in particular implies $Z$ does not admit a PSC metric.
  4. Proof of (4).
    By Proposition 5, the map $f:Z\to M$ is of positive degree. Moreover, by hypothesis $M$ admits a map of positive degree to an aspherical manifold. Composing these maps, we get a positive degree map from $Z$ to an aspherical manifold. Then by [CLL23, Corollary 3], $Z$ does not admit a PSC metric.
  5. Proof of (5).
    Since $M$ is symplectically aspherical, by Proposition 6, $M$ is a rationally essential manifold. Since the map $f:Z\to M$ if of positive degree by Proposition 5, $Z$ must be rationally essential as well. Now, by Lemma 2,
    $\pi_{1}(Z)\cong\pi_{1}(M\times S^{2})\cong\pi_{1}(M).$
    By hypothesis, $\pi_{1}(M)$ is amenable. Therefore, $Z$ is a rationally essential manifold with amenable fundamental group. Hence, by [Dra11], $\dim_{mc}Z=4$.
4.2. Proof of Corollary B
Before we move to the proof of Corollary B we give several examples of $4$-manifolds $M$ satisfying the hypotheses of Theorem A.
Example 8.
Let $X^{4}$ be a $T^{2}$-bundle over $T^{2}$. Then $M=X\#m\overline{\mathbf{CP}^{2}}$ satisfies the hypotheses required for Conclusions $(4)$ and $(5)$ of Theorem A.
Indeed, by a well-known observation of Thurston [Thu76], $X$ admits a symplectic structure. By the symplectic blowup construction due to Gompf [Gom95], $M$ also admits a symplectic structure. Note that $M$ is not aspherical for $m>1$ but admits a degree 1 map onto the aspherical manifold $T^{4}$. The fundamental group $\pi_{1}(X)$ is a $\mathbf{Z}^{2}$-extension over $\mathbf{Z}^{2}$, therefore it is amenable.
As a trivial case of the above, one could consider $X=T^{4}$. Note that here we also have $b_{2}^{+}=3>1$. Thus, $M$ satisfies the hypotheses required for Conclusions $(3),(4)$ and $(5)$.
Example 9.
Let $(X^{2n},\omega)$ be any symplectically aspherical manifold with amenable fundamental group. By repeated applications of Theorem 1 and Proposition 2, we can produce a symplectic submanifold $M^{4}\subset X^{2n}$ such that $\pi_{1}(M)\cong\pi_{1}(X)$. Since $X$ is symplectically aspherical, so is $M$. Therefore, $M$ satisfies the hypotheses required for Conclusion $(5)$ in Theorem A.
As an explicit example, one could take $X=T^{6}$. Considering the symplectic structure on $T^{6}$ as the product $T^{4}\times T^{2}$, we see that the projection $M\to T^{4}$ to the first factor must be a positive degree map by the same argument as in Proposition 5. Therefore, $M$ satisfies the hypotheses of Conclusions $(3),(4)$ and $(5)$. One salient feature of this example is that $\pi_{1}(M)\cong\mathbf{Z}^{6}$ has cohomological dimension $6$. Therefore, the techniques in [BD16] are unable to provide apriori bounds on the macroscopic dimension of $E_{Z}$.
Example 10.
(Examples within examples) Take any manifold $M$ satisfying the hypotheses required for Conclusions $(3),(4)$ and $(5)$ in Theorem A. Let $Z\subset M\times S^{2}$ be a Donaldson divisor as in the statement of Theorem A. Then $Z$ also satisfies the same hypotheses.
Proof of Corollary B.
Let $M$ be any manifold from Example 8, 9 or 10 above. Let $k\gg 1$ be sufficiently large. Let $Z_{k}\subset M\times S^{2}$ be the submanifold furnished by Donaldson’s theorem (Theorem 1), such that $[Z_{k}]=\mathrm{PD}(k\cdot[\omega])$. Let $L$ be the pre-quantum line bundle over $M$, so that $c_{1}(L)=[\omega]$. Let $E_{k}$ be the unit sphere bundle of $L^{\otimes k}$ restricted to $Z_{k}$. It follows from the proof of Theorem A that $Z_{k}$ and $E_{k}$ satisfy Conclusions $(1)$ and $(2)$. Proposition 4 implies $|\chi(Z_{k})|\to\infty$ as $k\to\infty$, as desired. ∎
4.3. Proof of Corollary C
We begin by recalling the definition of enlargeable manifolds from Gromov-Lawson [GL83].
Definition 4.
A compact orientable Riemannian manifold $M$ of dimension $n$ is called $\varepsilon^{-1}$-hyperspherical if there exists a $\varepsilon$-contracting map of positive degree from $M$ to the unit $n$-sphere. Further, $M$ is called enlargeable if for every $\varepsilon>0$, there exists a finite cover of $M$ whch is $\varepsilon^{-1}$-hyperspherical and spin.
The basic example of an enlargeable manifold is the unit circle $S^{1}\cong\mathbf{R}/2\pi\mathbf{Z}$. Indeed, the $n$-sheeted cover of $\mathbf{R}/2\pi\mathbf{Z}$ is $\mathbf{R}/2\pi n\mathbf{Z}$. The map $\mathbf{R}/2\pi\mathbf{Z}\to\mathbf{R}/2\pi n\mathbf{Z}$ given as scaling by $1/n$ is a $1/n$-contracting map, for any $n\geq 1$. Also, $S^{1}$ is certainly a spin manifold. It is straightforward to see that
  1. (1)
    Product of enlargeable manifolds is enlargeable,
  2. (2)
    A manifold $M$ admitting a positive degree map to an enlargeable manifold is also enlargeable, provided $M$ admits a finite spin covering.
Therefore, $T^{n}\cong S^{1}\times\cdots\times S^{1}$ is enlargeable and any spin manifold admitting a positive degree map to $T^{n}$ is also enlargeable. In [GL83], it is proved that all compact hyperbolic manifolds are enlargeable. The notion of enlargeability is important because of the main result of Gromov-Lawson [GL83, Theorem A], which shows that enlargeable manifolds do not admit PSC metrics.
Proof of Corollary C.
Let $M=T^{4}$ as in Example 8 with $m=0$. Let $\omega$ denote the product symplectic form on $T^{4}\times S^{2}$. Let $k\gg 1$ be a sufficiently large even integer. Then, by Theorem A, there exists a submanifold $Z^{4}\subset(T^{4}\times S^{2},\omega)$ such that $[Z]=\mathrm{PD}(k\cdot[\omega])$, $Z$ admits a positive degree map to $T^{4}$, and there exists a nontrivial circle bundle $E_{Z}\to Z$ such that the total space $E_{Z}$ admits a metric of positive scalar curvature.
Since $T^{4}$ is enlargeable and $Z$ already admits a positive degree map to $T^{4}$, to show that $Z$ is also enlargeable we just need to show that it is also spin, i.e. the second Stiefel-Whitney class is 00. Note that as $Z$ is a symplectic manifold, the second Stiefel-Whitney class of $Z$ is the first Chern class of $Z$ modulo $2$:
$w_{2}(Z)\equiv c_{1}(Z)\pmod{2}$
By Equation (2.3) in Proposition 4, we have
$c_{1}(Z)=j^{*}c_{1}(T^{4}\times S^{2})-k\cdot j^{*}[\omega]$
Next, we have $c_{1}(T^{4}\times S^{2})\equiv w_{1}(T^{4}\times S^{2})\equiv 0\pmod{2}$, as $T^{4}$ and $S^{2}$ are both spin, and product of spin manifolds is also spin. Since $k$ was moreover chosen to be even, we conclude $c_{1}(Z)\equiv 0\pmod{2}$. Thus, $w_{2}(Z)=0$. Hence, $Z$ is spin. ∎